OpenStudy (anonymous):

what is a 3x3 invertible matrix?

OpenStudy (tkhunny):

Are you OK with what a 3x3 matrix is?

OpenStudy (anonymous):

do you want a formal definition or a solved example showing all steps of inverting a 3x3 matrix?

OpenStudy (anonymous):

a solved example would be great

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@mathsmind ???

OpenStudy (anonymous):

which means its determinant cant be 0?

OpenStudy (anonymous):

yest i will show you the full steps

OpenStudy (anonymous):

consider the following matrix

OpenStudy (anonymous):

PLZ continue

OpenStudy (anonymous):

\[A=\left[\begin{matrix}1 & -2 & 0\\ 3& 1 & 5\\ -1 & 2 & 3\end{matrix}\right]\]

OpenStudy (anonymous):

Step one: \[A^T\]

OpenStudy (anonymous):

which means transpose of rows and columns of A

OpenStudy (anonymous):

\[A^T=\left[\begin{matrix}1 & 3 & -1\\ -2& 1 & 2\\ 0 & 5 & 3\end{matrix}\right]\]

OpenStudy (anonymous):

Step Two: find the adjoint or adjugate of A, which implies that we replace each element by their cofactor...

OpenStudy (anonymous):

If P = \begin{bmatrix} a &b &c \\ d &e &f \\ g &h &i \end{bmatrix}

OpenStudy (anonymous):

now please concentrate on this step it is simple but requires a bit of accuracy

OpenStudy (anonymous):

And P^4=2P how can we find P

OpenStudy (anonymous):

let's finish this first part then i will come to the 2nd question, you asked me first for a numerical example

OpenStudy (anonymous):

so we need to find adjA

OpenStudy (anonymous):

I know how to find a inverse matrix. So, can u plz do the 2nd question

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

yes sorry am on the phone, but just a quick hint for P^n we use the principle of induction

OpenStudy (anonymous):

i will show you shortly

OpenStudy (anonymous):

@mathsmind ???

OpenStudy (anonymous):

@mathsmind why is taking so long?

OpenStudy (anonymous):

sorry back, am at work

OpenStudy (anonymous):

now P^4 means you are multiplying your matrix by itself 4 times.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you know how to multiply the, it would be PP=P^2, then P^2P=P^3. then P^3P=P^4

OpenStudy (anonymous):

But it would be really long

OpenStudy (anonymous):

well this is the world of Matrices...

OpenStudy (anonymous):

are you thinking of algebraic operation on matrices such as taking the inverse of 2P

OpenStudy (anonymous):

or subtracting p^4-2P=0 in order to find the new P?

OpenStudy (anonymous):

Isn't there any short way

OpenStudy (anonymous):

have you done hermitian matrices...

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

listen can we kindly finish this tonight because it is 5 am and i need to go ...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i will provide you with full solution even if you are not online ok

OpenStudy (anonymous):

see ya tonight

OpenStudy (anonymous):

@ParthKohli

hero (hero):

lol, you still here? I thought you got help

OpenStudy (anonymous):

No

OpenStudy (anonymous):

can u do it?

OpenStudy (anonymous):

Short way

Parth (parthkohli):

I haven't done matrices yet: I am stuck with number theory at the moment.

Parth (parthkohli):

Nope.

OpenStudy (anonymous):

http://www.wikihow.com/Inverse-a-3X3-Matrix

OpenStudy (anonymous):

@looser there are many matrices such that P^4=2P but only one is invertible which is